OpenStudy (anonymous):

square root(x+10)+8=3

OpenStudy (anonymous):

square root has to be positive, isn't it?

OpenStudy (anonymous):

\[\sqrt{x+10} + 8 = 3\] \[\sqrt{x+10}=-5\] \[x + 10 = (-5)^{2}\] \[x+10=25\] \[x=15\]

OpenStudy (anonymous):

if you swapped the answer around the equal sign in the equation, both signs would change with it, making it true anyway